# A Closer Look at TabPFN v2: Strength, Limitation, and Extension

### v1

@@ 4 Comprehensive Evaluation of TabPFN v2
Building upon the remarkable performance of TabPFN v2, we conduct an extensive evaluation across diverse scenarios to assess its generalizability. Specifically, we evaluate TabPFN v2 on over 300 tabular datasets spanning a wide range of domains, scales, dimensionalities, and tasks.
Figure 2: Wilcoxon-Holm test at a significance level of 0.05 over 261 small- to medium-scale datasets. We omit the 12 datasets with more than 10 classes and the 27 datasets used to select the TabPFN v2 checkpoint from the 300 datasets in (Ye et al., 2024b).
@@ 4.1 Setups
To ensure a comprehensive evaluation, we adopt the benchmark proposed in Ye et al. (2024b), which includes 300 datasets covering 120 binary classification tasks, 80 multi-class classification tasks, and 100 regression tasks across various domains, dataset sizes, and feature types. This benchmark addresses common issues, such as mislabeled datasets and dataset redundancies due to multi-version overlaps (Kohli et al., 2024), ensuring high-quality evaluations.
Out of the 300 datasets, 12 classification datasets contain more than 10 classes, and 27 datasets overlap with the validation set used for checkpoint selection during TabPFN v2’s pre-training (Hollmann et al., 2025). To avoid evaluation bias, we exclude these datasets and report comparison results on the remaining 261 datasets.
Following the protocol in Gorishniy et al. (2021, 2024), each dataset is randomly split into training, validation, and test partitions in proportions to 64%/16%/20%. TabPFN v2 predicts the test set labels directly in an in-context learning fashion without further parameter and hyper-parameter tuning. Compared (deep) tabular methods perform hyper-parameter tuning using Optuna (Akiba et al., 2019) with 100 trials over the training set, and are early stopped on the validation set. All methods are evaluated with 15 random seeds, and the average performance across seeds is reported.
For classification tasks, we use accuracy (higher is better) as the evaluation metric, while regression tasks are evaluated using Root Mean Square Error (RMSE, lower is better).
Figure 3: Probability of Achieving the Best Accuracy/RMSE (PAMA) across 261 small- to medium-scale datasets. The values inside the rectangles represent the percentage of datasets where each method achieves the best accuracy.
4.2 Results
We compare TabPFN v2 against 26 representative (deep) tabular methods on 261 datasets (detailed references are in the appendix). To assess statistical significance, we apply the Wilcoxon-Holm test at a 0.05 significance level (Demsar, 2006). The critical difference analysis results are presented in Figure 2.
The results demonstrate that TabPFN v2 consistently outperforms both tree-based methods, such as CatBoost (Prokhorenkova et al., 2018), and deep tabular models, including RealMLP (Holzmüller et al., 2024), ModernNCA (Ye et al., 2025), TabR (Gorishniy et al., 2024), and FT-Transformer (Gorishniy et al., 2021), with statistically significant differences.
To further analyze performance, we consider the Probability of Achieving the Best Accuracy/RMSE (PAMA) (Delgado et al., 2014), which measures the percentage of datasets where a method achieves the best performance. As shown in Figure 3, TabPFN v2 achieves the highest PAMA score, with 25.06% of datasets yielding the best results, outperforming other methods like ModernNCA (with a percentage of 13.73%) and CatBoost (9.48%). These results validate TabPFN v2’s exceptional generalizability on small- to medium-scale tabular datasets, further enhanced by its efficiency and lack of reliance on hyper-parameter tuning.
Table 1: Average rank (lower is better) of TabPFN v2 and representative comparison methods on 18 high-dimensional datasets and 18 large-scale datasets. Full results with our extensions are in Figure 6 and Figure 8.
$\downarrow$ TabPFN v2 CatBoost MNCA RealMLP LR
@@ 4.2 Results
4.3 Additional Evaluations
@@
To examine TabPFN v2’s performance on larger and more complex datasets, we conduct additional evaluations on 18 large-scale datasets with $N\times d>1,000,000$ and 18 high-dimensional datasets with $d\geq 2,000$ (Jiang et al., 2024). For high-dimensional datasets, we follow the same evaluation protocol as before. For large-scale datasets, due to the prohibitive cost of hyper-parameter tuning, default hyper-parameters are used for all methods. The average ranks of some representative methods are summarized in Table 1. Full results are in section 7 with our extensions.
While TabPFN v2 performs superiorly on small- to medium-scale datasets, its performance degrades on large-scale and high-dimensional datasets. For instance, on large-scale datasets, TabPFN v2 ranks lower than both CatBoost and RealMLP, highlighting its scalability limitations. On high-dimensional datasets, TabPFN v2 fails to outperform even the simple Logistic Regression (LR) model, further emphasizing the need for enhancements to better handle these challenging scenarios.
These results highlight TabPFN v2’s limitations in handling high-dimensional and large-scale datasets, which may result from two key factors. First, TabPFN v2 is pre-trained exclusively on small- and medium-scale synthetic datasets, creating a mismatch when applied to large-scale datasets. Second, the complexity of transformers in TabPFN v2 scales with both $N$ (the number of instances) and $d$ (the number of features), meaning that increasing either value significantly can lead to a substantial rise in computational complexity, which in turn may reduce the model’s effectiveness.
In summary, while TabPFN v2 demonstrates exceptional performance on small- to medium-scale datasets, scaling it up to handle larger datasets, higher-dimensional feature spaces, and multi-class tasks exceeding 10 classes is a crucial and promising direction for future research.
5 Heterogeneity Mechanisms Analysis
As described in section 3, TabPFN v2 handles data heterogeneity by embedding a $d$-dimensional input vector into a $k$-dimensional space with size $d\times k$. This enables TabPFN v2 to adapt to diverse downstream datasets with varying feature dimensions seamlessly. In this section, we analyze this mechanism and revisit it as a variant of the token-based deep tabular method. Further analyses of the key components in TabPFN v2 are provided in Appendix D.
5.1 Analysis of the Heterogeneity Handling Strategy
Traditional deep tabular models often struggle with dataset heterogeneity due to their reliance on fixed feature semantics. For example, token-based approaches like AutoInt (Song et al., 2019) and FT-Transformer (FT-T) (Gorishniy et al., 2021) learn dataset- and attribute-specific feature tokens during training, which become ineffective when applied to new datasets with different feature meanings. These methods typically generate $d$ attribute-specific tokens $[{\bm{r}}_{1},\ldots,{\bm{r}}_{d}]\in{\mathbb{R}}^{k\times d}$, one for each dimension, and transform an instance ${\bm{x}}_{i}$ into $[{x_{i}^{1}}\cdot{\bm{r}}_{1},\ldots,{x_{i}^{d}}\cdot{\bm{r}}_{d}]\in{\mathbb{R}}^{k\times d}$, where $x_{i}^{j}$ is the $j$-th element of ${\bm{x}}_{i}$. However, when transitioning to a new dataset with $d^{\prime}$ features, these learned tokens cannot be reused, requiring feature tokens to be learned anew. Previous methods address this by relying on semantic descriptions of attributes (Yan et al., 2024) or assuming overlap between feature sets to construct a mapping (Onishi et al., 2023; Zhou et al., 2023).
In contrast, TabPFN v2 eliminates this dependency through randomized feature tokens. Specifically, TabPFN v2 applies a shared $k$-dimensional vector ${\bm{u}}\in\mathbb{R}^{k}$ to lift raw attribute values into a $k$-dimensional space. To differentiate attributes, TabPFN v2 adds a random $k$-dimensional perturbation for each attribute, consistent within a dataset but distinct across datasets. Formally, the randomized tokens are denoted as:
${\bm{R}}={\bm{W}}{\bm{P}}\in{\mathbb{R}}^{k\times d},$ (1)
where ${\bm{P}}\in{\mathbb{R}}^{k^{\prime}\times d}$ is randomly generated, and ${\bm{W}}\in{\mathbb{R}}^{k\times k^{\prime}}$ is a learnable pre-condition projection matrix that maps the random values to the required dimensionality ($k^{\prime}<k$).The $j$-th column of ${\bm{R}}$, ${\bm{r}}_{j}\in{\mathbb{R}}^{k}$, serves as the specific token for the $j$-th attribute and is combined with the shared vector ${\bm{u}}$. Then, an instance ${\bm{x}}_{i}$ is transformed through
$[{x_{i}^{1}}\cdot({\bm{u}}+{\bm{r}}_{1}),\ldots,{x_{i}^{d}}\cdot({\bm{u}}+{\bm{r}}_{d}),\tilde{{\bm{y}}}_{i}]\in{\mathbb{R}}^{k\times(d+1)}\;.$ (2)
This randomized tokenization ensures that distinct attributes—irrespective of their semantic meanings—are mapped to unique, nearly orthogonal directions in the high-dimensional space. This is akin to one-hot encoding but operates in a continuous over-parameterized regime, where random vectors naturally approximate orthogonality.
Through its transformer layers, TabPFN v2 dynamically combines these basis vectors during inference, weighting their contributions based on the input context. This process is analogous to inverse PCA: while PCA decomposes data into orthogonal components, TabPFN v2 reconstructs heterogeneous features by mixing randomized bases scaled by their attribute values. This approach avoids the need for pre-defined feature semantics, while the shared vector ${\bm{u}}$ introduces a stable inductive bias that supports cross-dataset generalization. Pre-training on diverse and abundant synthetic datasets further equips TabPFN v2 with the capability to adapt these bases across diverse tasks, unifying disparate tabular datasets into a coherent latent framework.
This randomized token strategy effectively removes the need for dataset- and attribute-specific feature token learning, addressing a significant limitation of prior methods and improving transferability to diverse downstream datasets.
Figure 4: The change of ranks between FT-T, MNCA, and the variants with randomized tokens over 45 tabular datasets. (The rank is calculated together with the 26 compared baselines.) Values in the parentheses denote the average rank.
5.2 Effectiveness of Randomized Feature Tokens
To further evaluate the effectiveness of randomized tokens, we integrate this strategy with two representative methods: FT-Transformer (FT-T) by Gorishniy et al. (2021) and ModernNCA (MNCA) by Ye et al. (2025).
FT-T with Randomized Tokens (FT-T). FT-T is a token-based method that learns specific tokens $\{{\bm{r}}_{1},\ldots,{\bm{r}}_{d}\}\in{\mathbb{R}}^{d\times k}$ for each attribute. After transforming an instance into a set of tokens, a transformer processes these tokens together with a learned class token. The output token corresponding to the class token is then fed into a linear classifier.
We replace FT-T’s learned tokens with the randomized tokens used in TabPFN v2, retrain FT-T (denoted as FT-T), and evaluate the resulting performance.
@@ 5.2 Effectiveness of Randomized Feature Tokens
Figure 5: Visualization of extracted embeddings for four datasets: churn (first row, two classes), bank (second row, two classes), KDD (third row, two classes), and website_phishing (fourth row, three classes). We use crosses to denote training examples and circles to denote test examples. (a) shows the raw features, while (b) presents the embeddings extracted using the vanilla strategy. (c)-(f) depict the embeddings obtained using our proposed methods across different layers. The accuracy value is calculated by training a linear model (logistic regression) over the extracted embeddings on the training set and predicting on the test set.
MNCA with Randomized Tokens (MNCA). ModernNCA is a neighborhood-based method that predicts an instance based on its similarity to training instances in a latent space. While ModernNCA typically operates on raw features, we adapt it into a token-based method by applying ModernNCA to the token output from FT-T before the final linear layer, incorporating the randomized tokens as above.
We evaluate these FT-T and MNCA variants over the tiny-bechmark2 from Ye et al. (2024b), which contains 45 representative tabular datasets. The results in Figure 4 demonstrate how the performance rank changes when replacing the specific tokens with the randomized tokens.
Remark. Our experiments confirm that feature tokens do not significantly rely on semantic meanings. Using randomized tokens only slightly increases the average ranks as long as they are distinct across attributes. This highlights the robustness and universality of the randomized token strategy, demonstrating its effectiveness in handling heterogeneous datasets and its potential to improve generalization across diverse tasks once pre-trained.
6 TabPFN v2 as a Feature Encoder
To better understand TabPFN v2, we examine whether its in-context learning capability also produces separable and meaningful feature representations.
6.1 Challenges in Embedding Extraction
As described in section 3, the output tokens from the multiple transformer layers in TabPFN v2 correspond one-to-one with the input tokens, forming a tensor of size $(N+1)\times(d+1)\times k$. The final token, corresponding to the (dummy) label embedding $\tilde{{\bm{y}}}^{*}$ of the test instance, is passed through an MLP block to produce the output. An intuitive idea is to treat the output tokens associated with the training label embeddings $\{{\bm{y}}_{i}\}_{i=1}^{N}$—prior to the MLP block—as the extracted embeddings for the training data.
However, this straightforward method has one fundamental limitation due to the distinct roles of labeled training and unlabeled test data in TabPFN v2’s in-context learning process. Specifically, the label embeddings for the training instances are derived from their true labels, while those for the test instances rely on dummy labels. This discrepancy results in output embeddings that are inherently non-comparable between the training and test instances.

### v2

@@ 4 Comprehensive Evaluation of TabPFN v2
Before presenting our core studies, we first extend TabPFN v2’s evaluation beyond the original set of datasets to over 300, covering a much broader range of domains, attributes, scales, dimensionalities, and tasks [24, 50, 76, 61], aiming to more thoroughly assess its generalizability and limitations.
@@ 4.1 Setups
We first adopt the benchmark from [76], comprising 120 binary classification, 80 multi-class classification, and 100 regression tasks. It resolves common issues such as mislabeled data and redundancies from overlapping dataset versions [43], enabling more reliable evaluations. Out of the 300 datasets, 27 belong to the validation set used for checkpoint selection in TabPFN v2 [29]. To avoid evaluation bias, we exclude these datasets and report results on the remaining 273 datasets.
Following the protocol in [22, 23], each dataset is randomly split into training, validation, and test partitions in a 64%/16%/20% ratio. TabPFN v2 predicts test set labels directly using in-context learning, without any additional parameter or hyperparameter tuning. Baseline tabular methods—both deep and traditional—perform hyperparameter tuning using Optuna [1], with 100 trials on the training set and early stopping based on validation performance.
All methods are evaluated using 15 random seeds, and we report the average performance across seeds. For classification tasks, we report accuracy (higher is better), while regression tasks are evaluated using Root Mean Square Error (RMSE; lower is better). For tasks with more than 10 classes, we adopt the built-in Error-Correcting Output Codes (ECOC) [16] strategy for TabPFN v2.
Figure 2: Probability of Achieving the Maximum Accuracy or Minimum RMSE across 273 datasets. Values inside rectangles show the percentage of datasets on which a method achieves the best result.
$\downarrow$ TabPFN v2 CatB MNCA R-MLP LR
@@
>10 classes 3.33 2.75 3.17 1.42 4.33
Figure 3: Average rank (lower is better) of TabPFN v2 and representative baselines on 18 high-dimensional, 18 large-scale, and 12 datasets with more than 10 classes. Full results with our extensions are in Figure 6.
4.2 Empirical Results: Strengths of TabPFN v2
We compare TabPFN v2 against 26 representative tabular methods (see Appendix A for full references). To assess statistical significance, we apply the Wilcoxon-Holm test with a significance level of 0.05 [14]. As shown in the critical difference diagram in Figure 1 (right), TabPFN v2 consistently outperforms both tree-based methods, such as CatBoost [57], and deep tabular models, including RealMLP [30], ModernNCA [77], TabM [20], TabR [22], and FT-Transformer [23].
To further assess performance, we report the Probability of Achieving the Maximum Accuracy or Minimum RMSE (PAMA) [13], which measures the proportion of datasets on which a method achieves the best performance. As shown in Figure 3, TabPFN v2 attains the highest score, delivering the top results on 26.02% of the datasets—outperforming other methods such as ModernNCA (11.79%) and TabM (9.78%). These results underscore TabPFN v2’s strong generalizability.
4.3 Empirical Results: Limitations of TabPFN v2
The above evaluation focuses on small- to medium-scale datasets, specifically those with fewer than 10,000 examples. However, as noted in [29], the computational complexity of transformers constrains TabPFN v2’s ability to scale effectively to datasets with larger sample sizes or higher dimensionality.
@@ 4.3 Empirical Results: Limitations of TabPFN v2
To verify this, we conduct additional evaluations on 18 high-dimensional datasets with $d\geq 2{,}000$ [37] and 18 large-scale datasets where $N\times d>1{,}000{,}000$. For high-dimensional datasets, we follow the same protocol as before. For large-scale datasets, due to the prohibitive cost of hyperparameter tuning, default hyperparameters are used for all methods. The average ranks of several representative methods are summarized in Figure 3. The full results—along with our extensions—are in Section 7.
As shown, TabPFN v2’s performance degrades on both large-scale and high-dimensional datasets. On large-scale datasets, it ranks below both CatBoost and RealMLP; on high-dimensional datasets, it even falls behind the simple Logistic Regression (LR) model. Beyond these two limitations, Figure 3 also reports results on the 12 datasets in Section 4.2 that contain more than 10 categories, where the ECOC strategy currently used by TabPFN v2 appears ineffective in achieving high accuracy. While increased computational complexity may contribute to this reduced effectiveness, we hypothesize that TabPFN v2 was pre-trained exclusively on small- to medium-scale synthetic datasets with fewer than 10 categories, leading to a mismatch when applied to larger or more complex real-world data.
These results underscore the limitations of TabPFN v2, suggesting areas for further improvement.
5 How Does TabPFN v2 Effectively Handle Data Heterogeneity?
Section 4 demonstrates TabPFN v2’s excellent generalizability to heterogeneous downstream tasks while also highlighting its current limitations. In the rest of the paper, we first examine the mechanisms behind its strengths, followed by methods to overcome its limitations.
5.1 Diving into TabPFN v2’s Mechanisms for Heterogeneous Input
Revisiting the problem. As noted in Sections 3 and 2, tabular datasets often differ in both the number of attributes (i.e., $d$) and the semantics of those attributes. Even when dimensionalities match, the dimensional semantics from different datasets are typically not directly comparable. A robust tabular foundation model must therefore handle such heterogeneity effectively, enabling it to learn from diverse pre-training datasets and transfer its capabilities to new downstream tasks.
Tokenization as a feasible solution. Among prior approaches, the most relevant to TabPFN v2 are token-based methods [65, 23, 75]. The core idea is to convert a $d$-dimensional instance ${\bm{x}}\in\mathbb{R}^{d}$ into a set of $d$ fixed-dimensional tokens (each of dimension $k$), with one token per attribute. This enables the use of transformer architectures, which naturally accommodate variability in $d$ across datasets.
To embed each attribute into a shared $k$-dimensional space, prior work either uses pre-defined semantic embeddings [75] (e.g., word vectors of attribute names) or learns dataset-specific embeddings [65, 23]. Given $d$ attribute-specific tokens $[{\bm{r}}_{1},\ldots,{\bm{r}}_{d}]\in{\mathbb{R}}^{d\times k}$, each instance ${\bm{x}}_{i}\in{\mathbb{R}}^{d}$ can then be transformed into $\left[x_{i}^{1}\cdot{\bm{r}}_{1},\ldots,x_{i}^{d}\cdot{\bm{r}}_{d}\right]\in{\mathbb{R}}^{d\times k}$, where $x_{i}^{j}$ denotes the $j$-th element of ${\bm{x}}_{i}$.
By embedding all attributes into a shared, fixed-dimensional feature space, this approach allows the transformer to learn transferable patterns and knowledge from heterogeneous datasets.
Difficulty in direct generalization. While appealing, the aforementioned methods face a notable challenge when applied to downstream tasks: attribute names or semantics are not always accessible, as discussed in Section 2. Although it is possible to learn dataset-specific attribute tokens, doing so incurs additional computational cost and prohibits the reuse of previously learned tokens. Consequently, this limits the direct generalization of the foundation model to new tasks.
TabPFN v2’s mechanisms. TabPFN v2 builds on prior token-based methods by representing each instance ${\bm{x}}_{i}$ as a sequence of tokens. However, rather than assigning a deterministic token to each attribute, TabPFN v2 samples random tokens at inference time. Specifically, it learns a shared vector ${\bm{u}}\in\mathbb{R}^{k}$ that lifts each element of ${\bm{x}}_{i}$ into a $k$-dimensional space. To distinguish attributes, TabPFN v2 adds a random perturbation to each one. [^2] This detail was identified from the supplementary material and code of [29]. For the $j$-th attribute (i.e., $x_{i}^{j}$), the representation becomes $x_{i}^{j}\cdot{\bm{u}}+{\bm{r}}_{j}$, where ${\bm{r}}_{j}={\bm{W}}{\bm{p}}_{j}$. Here, ${\bm{p}}_{j}\in\mathbb{R}^{k^{\prime}}$ is a randomly generated vector, and ${\bm{W}}\in\mathbb{R}^{k\times k^{\prime}}$ is a learned projection matrix that conditions the perturbation. The full instance ${\bm{x}}_{i}$ is then represented as:
$[{x_{i}^{1}}\cdot{\bm{u}}+{\bm{r}}_{1},\ldots,{x_{i}^{d}}\cdot{\bm{u}}+{\bm{r}}_{d},\tilde{{\bm{y}}}_{i}]\in{\mathbb{R}}^{k\times(d+1)},$ (1)
where the last token $\tilde{{\bm{y}}}_{i}$ encodes the label information (see Figure 1 for an illustration).
(b) Layer 3
(c) Layer 6
(d) Layer 9
(e) Layer 12
(f) Legend
Figure 4: Attribute relationships inferred by TabPFN v2. The first and third rows show PCA projections of the $d$ attribute tokens from all $N$ training instances at various layers for the churn and bank datasets. Colors indicate different attributes (see legend on the right). The second and fourth rows display the attribute-wise attention maps. Each matrix cell represents the average attention weight between attributes; the last element along each axis (e.g., the last column and row) corresponds to the label. The first plots in the second and fourth rows summarize the cosine similarity of attention maps across random seeds. See text for details. The abbreviations of feature names of these two datasets are explained in Table 4 and Table 5.
5.2 TabPFN v2 Internalizes Attribute Token Learning
TabPFN v2’s randomized tokenization scheme eliminates the need to define attribute- or dataset-specific tokens across tasks, thereby syntactically enabling direct application of the pre-trained model. At first glance, this may appear to disregard the valuable semantic meaning of attributes. However, we show that through in-context learning, TabPFN v2 can consistently infer relationships among attributes within a dataset—despite the randomness introduced during tokenization. Specifically, we analyze the behavior of attribute tokens from three perspectives, as illustrated in Figure 4, using two representative downstream datasets: churn and bank.
First, we visualize the attribute token embeddings (i.e., the first $d$ tokens in Equation 1) across all $N$ training instances. The first and third rows of Figure 4 present PCA projections of these $N\times d$ tokens at the input stage and after transformer layers $\{3,6,9,12\}$, with colors indicating different attributes. Initially, tokens from different attributes appear randomly scattered. However, as the input progresses through the transformer layers, these tokens become increasingly structured. For example, in the bank dataset (which predicts term deposit subscriptions), attributes such as “job,” “education,” and “balance” eventually cluster into semantically coherent groups.
Second, we examine attribute-wise attention patterns across layers, including attention to the label token. The second and fourth rows of Figure 4 show heatmaps of attention weights averaged over heads and training instances. Each row in the heatmap represents the attention distribution from one attribute to all others; the last element along each axis (e.g., the last column and row) corresponds to the label. Darker shades indicate stronger attention. We observe a consistent pattern across datasets: in early layers, attributes predominantly attend to the label token, likely to absorb task-specific signals. In intermediate layers, attention becomes more uniformly distributed, facilitating information exchange across attributes. In deeper layers, attention concentrates on semantically relevant attributes, suggesting the model has inferred inter-attribute relationships useful for prediction.
Lastly, to assess robustness against random token initialization, we compute attribute-wise attention weights across 10 runs. The cosine similarities and variances of these attention patterns are summarized in the first plots of the second and fourth rows in Figure 4. The results confirm that attention patterns remain stable across runs, except for the first layer.
Remark. The above results suggest that TabPFN v2 can reliably infer meaningful attribute relationships through in-context learning. Although input embeddings are randomized, they consistently differentiate attributes across instances—functionally akin to one-hot encodings. Pre-training on diverse tasks thus enables the model to extract predictive patterns (e.g., co-occurrence across attributes, value distributions, and relative magnitudes) directly from the statistical structure of each dataset, without relying on pre-defined attribute semantics. As a result, the model effectively internalizes attribute token learning within the inference process.
6 TabPFN v2 Can Be Transformed into an Effective Feature Encoder
In Section 5, we show that TabPFN v2’s in-context learning process infers meaningful attribute relationships. Here, we examine whether TabPFN v2 also produces separable instance representations.
@@ 6 TabPFN v2 Can Be Transformed into an Effective Feature Encoder
Figure 5: Visualization of the extracted instance features from two datasets: churn (first row, binary) and bank (second row, binary). Blue and red indicate classes; darker crosses and lighter circles denote training and test samples. (a) shows the raw input features (e.g., ${\bm{x}}_{i}$), while (b) presents embeddings from the vanilla strategy. (c)-(f) display embeddings produced by our method at different layers. Classification accuracy is reported by training a linear logistic regression model on the training embeddings and evaluating on the test set.
6.1 Naive Feature Extraction Fails
As shown in Figure 1 (left), TabPFN v2 makes predictions based on the output token corresponding to the (dummy) label embedding $\tilde{{\bm{y}}}^{*}$ of the test instance. This output token can thus be interpreted as the instance embedding for the test example. A natural extension to obtain embeddings for the training instances is to extract the output tokens corresponding to the training label embeddings $\{\tilde{{\bm{y}}}_{i}\}_{i=1}^{N}$.
However, as shown in Figure 5 (b), this naive approach leads to surprisingly discrepant feature distributions between training (darker cross) and test (lighter circle) examples. As a result, a linear classifier trained on these embeddings performs poorly on the test set. We attribute this discrepancy to the distinct roles of labeled training data and unlabeled test data in TabPFN v2’s in-context learning process. Specifically, the label embeddings for the training instances are derived from true labels, whereas those for the test instances rely on dummy labels. This mismatch renders the resulting output embeddings non-comparable between training and test instances.